Claudia has test scores of 78, 88, 84, 92, 76, 98. If she wants to end up with a ''B'' average, what does she need to score on her next test?
Answer: she need to get an 86 to get a b average: 80 on her next test
Is this correct, my answer is 86.

Answer by LinnW(1048)

(78+ 88+ 84+ 92+ 76+ 98+ x)/7 = 86
(516+ x)/7 = 86
multiply each side by 7
516+ x = 602
add -516 to each side
x = 86
If 86 is a B, then she needs an 86 on her next test
to have an 86 or B average.